Page tree
Skip to end of metadata
Go to start of metadata

Attendees

Apologies

Agenda

      • VM status - Engineering.
      • VM product status.
      • AOB

Discussion items

ItemNotes
Introduction

RC apologized for not having minuted the previous meeting yet.

VM Status - engineering.

EUI:
Emil is still close to completing a new VM version but has not managed to return to that work yet.

EPD:

(not represented) AW has now had confirmed that the STEP update will not change the LL products.

SWA:

(not represented) AW mentioned that a version fixing the HIS files is still outstanding.

PHI:

MK and RC described the short series of VM deliveries leading to 01.08.07. MK will send RC a link/links to descriptions of the general issue with this kernel configuration and VM ware, in case useful for future reference.

MK described that the new version, 1) will report the correct CREATOR and VERS_SW in the headers, and 2) has an updated converter. He was not aware of the specific contents of that update.

SoloHI:

00.01.28 has been deployed since July in operations.
ADG will look at the files and send flight output to Lynn for her to examine.

Metis

DW (before arrival of Maurizio): Output files (from VM repackaged by RC) are much better. Close to workable. A few fixes still needed.

Maurizio (remarks actually made in AoB) understands the problems in the products. Does not yet understand the VM deployment problems. He will try to follow the procedure described by RC which managed to make the ova usable at ESAC. RC mentioned that he hoped more that (medium-term at least) it would lead to understanding the root of the problem.

SPICE:

(not represented) DW mentioned that we have a new verion from August. But that the changes are in keywords which are transparent to the SOC.

MAG

Ed not present. AW mentioned that the ops calibration was updated yesterday and seemed to have worked with data from the latest pass.

RPW:

Not represented.

STIX

Shane had understood why the VM was not processing the most recent TM. But this has been very hard to fix. He is on the point of rebuilding the VM from scratch.
He asked if there was an update to the SEGU re the interface between the VM and the SOC environment. RC said he thought that was stable now for a long time, but Shane should say which SEGU version he has and RC would provide the latest (NB there are change records in the document).

Shane is not working on this task yet; he is currently on fits issues for the SOAR. The VM will be his next task unless he finds someone else to do it.

DW asked if Shane yet has a description of the science products. Shane said he wants first to fix the fits files and then will move on to the document.

VM product status

Nothing to add, after a few points mentioned during above topic.

AoB

KS asked if, since the agendas from these meeting are getting thin, whether we should arrange them on-demand (as the agenda fills) rather than regularly.

We agreed that this did not have to change immediately, and that the SOC should consider the issue and make a proposal.

We agreed that there will be a next telecon and that it will be on the 1st of December.

Preparatory Actions

None.

New Actions

  • SOC to propose regularity for future meetings.